OpenStudy (laronjaxon):

x2-4x+4/x2+10+25*x+5/x2+3x-10 multipluying and dividing rational expressions

OpenStudy (anonymous):

please recheck the question... is that true x2+10+25?

OpenStudy (laronjaxon):

10x* sorry

OpenStudy (anonymous):

oh, ok..

OpenStudy (campbell_st):

well fractorising may help solve some difficulties \[\frac{(x - 2)^2}{x + 5)^2} \times \frac{ (x + 5)}{(x + 5)(x - 2)}\] that makes it easier... cancel some common factors and off you go

OpenStudy (anonymous):

\[\frac{ x ^{2}-4x+4 }{ x ^{2}+10x+25 }\times \frac{ x+5 }{ x ^{2}+3x-10 }\]\[\frac{ (x-2)(x-2) }{ (x+5)(x+5) }\times \frac{ (x+5) }{ (x+5)(x-2) }\]\[\frac{ x-2 }{ (x+5)(x+5) }\]\[\frac{ x-2 }{ x ^{2}+10x+25 }\]

OpenStudy (campbell_st):

@EchaSetiaWijaya the aim of the site is to help people get an answer... not provide the answer...

OpenStudy (anonymous):

yess, but i can't stand to solve the question. sorry

OpenStudy (laronjaxon):

Thank you @EchaSetiaWijaya